OpenStudy (anonymous):

The longer leg of a 30°-60°-90° triangle is 18. What is the length of the other leg?

OpenStudy (anonymous):

18 root 3

OpenStudy (anonymous):

@Devanshi The answers are 9 root 3, 12 root 3, 6 root 3 and 9

OpenStudy (dumbcow):

 this as ratio of 30-60-90 triangle set up proportion \[\frac{x}{1} = \frac{18}{\sqrt{3}}\]